# Hierarchical Position Embedding of Graphs with Landmarks and Clustering for Link Prediction

### v1

@@ Abstract.
Learning positional information of nodes in a graph is important for link prediction tasks. We propose a representation of positional information using representative nodes called landmarks. A small number of nodes with high degree centrality are selected as landmarks, which serve as reference points for the nodes’ positions. We justify this selection strategy for well-known random graph models and derive closed-form bounds on the average path lengths involving landmarks. In a model for power-law graphs, we prove that landmarks provide asymptotically exact information on inter-node distances. We apply theoretical insights to practical networks and propose Hierarchical Position embedding with Landmarks and Clustering (HPLC). HPLC combines landmark selection and graph clustering, where the graph is partitioned into densely connected clusters in which nodes with the highest degree are selected as landmarks. HPLC leverages the positional information of nodes based on landmarks at various levels of hierarchy such as nodes’ distances to landmarks, inter-landmark distances and hierarchical grouping of clusters. Experiments show that HPLC achieves state-of-the-art performances of link prediction on various datasets in terms of HIT@K, MRR, and AUC. The code is available at https://github.com/kmswin1/HPLC.
@@ 1. Introduction
Message passing GNNs (MPGNNs) (Gilmer et al., 2017; Kipf and Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018) have been successful in learning structural node representations through neighborhood aggregation. However, MPGNNs achieved subpar performances on link prediction tasks due to their inability to distinguish isomorphic nodes. Subgraph-based methods (Zhang and Chen, 2018; Pan et al., 2022) were proposed for link prediction based on dynamic node embeddings from enclosing subgraphs to tackle isomorphism. Labeling methods (Li et al., 2020; Zhang et al., 2021) assigns node labels to identify each node via hashing or aggregation functions. An important line of research is the position-based methods  (You et al., 2019; Dwivedi et al., 2020; Kreuzer et al., 2021; Wang et al., 2022) which propose to learn and utilize the positional information of nodes.
@@
The positional information of nodes is useful for link prediction because, e.g., isomorphic nodes can be distinguished by their positions, and the link formation between a node pair may be related to their relative distance (Srinivasan and Ribeiro, 2020). The position of a node can be defined as its distances relative to other nodes. A relevant result by Bourgain (Bourgain, 1985) states that inter-node distances can be embedded into Euclidean space with small errors. Linial (Linial et al., 1995) and P-GNN (You et al., 2019) focus on constructing Bourgain’s embeddings based on the nodes’ distances to random node sets. Another line of positional encoding used the eigenvectors of graph Laplacian (Dwivedi and Bresson, 2021; Kreuzer et al., 2021) based on spectral graph theory. However, for large or dense graphs, the aforementioned methods have convergence issues and may not scale well (Wang et al., 2022). For example, Laplacian methods have stability issues (Wang et al., 2022) and may not outperform methods based on structural features (Zhang and Chen, 2018). Thus, there is significant room for improving both performance and scalability of positional node embedding for link prediction.
@@
In this paper, we propose an effective and efficient representation of the nodes’ positional information. We select a small number of representative nodes called landmarks and impose hierarchy on the graph by associating each node with a landmark. Each node computes distances to the landmarks, and each landmark computes distances to other landmarks. Such distance information is combined so as to represent nodes’ positional information. Importantly, we select landmarks and organize the graph in a principled way, unlike previous methods using random selection, e.g., (Bourgain, 1985; Linial et al., 1995; You et al., 2019).
@@
The key question is how to select landmarks. We advocate the selection based on degree centrality which is motivated from the theory of network science. In network models with preferential attachment (PA), nodes with high degrees called hubs are likely to emerge (Barabási and Albert, 1999). PA is a process such that, if a new node joins the graph, it is more likely to connect to nodes with higher degrees. Node degrees follow power-law distribution, and the network exhibits scale-free property (Barabási and Albert, 1999). Hubs are abundant in social/citation networks and World Wide Web, and play a central role in characterizing the network.
@@
Motivated by the theory, we propose Hierarchical Position embedding with Landmarks and Clustering (HPLC). HPLC partitions a graph into $O(\log N)$ clusters which are locally dense, and appoints the node with the highest degree in the cluster as the landmark. Our intention is to bridge gap between theory and practice: hubs may not be present in real-world graphs. Thus, it is important to distribute the landmarks evenly over the graph so that nodes can access nearby/local landmarks. Next, we form a graph of higher hierarchy, i.e., the graph of landmarks, and compute its Laplacian encoding based on the inter-landmark distances. The encoding is assigned as a membership to the nodes belonging to the cluster, so as to learn positional features at the cluster level. We further optimize our model using the encoding based on hierarchical grouping of clusters. The computation of HPLC can be mainly done during preprocessing, incurring low computational costs. Our experiments on 7 datasets with 16 baseline methods show that HPLC achieves state-of-the-art performances in most cases, demonstrating its effectiveness over prior position-/distance-based methods.
Our contributions are summarized as: 1) we propose HPLC, a highly efficient algorithm for link prediction using hierarchical position embedding based on landmarks combined with graph clustering; 2) building upon network science, we derive closed-form bounds on average path lengths via detours for well-known random graphs which, to our belief, are important theoretical findings; 3) we conduct extensive experiments to show that HPLC achieves state-of-the-art performances in most cases.
@@ 2.1. Notation
We consider undirected graph $G=(V,E)$ where $V$ and $E\subseteq V\times V$ denote the set of vertices and edges, respectively. Let $N$ denote the number of nodes, or $N=|V|$. $\mathbf{A}\in\mathbb{R}^{N\times N}$ denotes the adjacency matrix of $G$. $d(v,u)$ denotes the geodesic (shortest-path) distance between $v$ and $u$. Node attributes are defined as $X=\{x_{1},...,x_{N}\}$ where $x_{i}\in\mathbb{R}^{n}$ denotes the feature vector of node $i$. We consider methods of embedding $X$ into latent space as $Z=\{z_{1},...,z_{N}\},z_{i}\in\mathbb{R}^{l}$. We study the node-pair-level task of estimating the link probability between $u$ and $v$ from $z_{u}$ and $z_{v}$.
@@ 2.2. Representation of Positions using Landmarks
$D(v)$ can be used to represent the (approximate) position of node $v$. If the positions of $u$ and $v$ are to be used for link prediction, one can combine $D(u)$ and $D(v)$ (or the embeddings thereof) for estimating the link probability.
@@
We examine the achievable accuracy of distance estimates based on the detours via landmarks. Previous estimates using auxiliary nodes relied on heuristics (Potamias et al., 2009) or complex algorithms for arbitrary graphs (Linial et al., 1995; You et al., 2019). However, real-world graphs have structure, e.g., preferential attachment or certain distribution of node degrees. Presumably, such structure can be exploited to devise efficient algorithms. The key design questions are: how to select good landmarks, and how many of them? Drawing upon the theory of network science, we analyze a well-known class of random graphs, derive the detour distances via landmarks, and glean design insights from the analysis.
(1) $P(L_{ij}>s)=\exp\bigg{[}-\frac{h_{i}h_{j}}{\beta N\langle h^{2}\rangle}\cdot\langle h^{2}\rangle_{Q}\cdot K(N)\cdot(s-1)\bigg{(}\frac{\langle h^{2}\rangle N}{\beta}\bigg{)}^{s-1}\bigg{]},\quad s=1,2,\cdots$ (2) $\bar{l}\leq\dfrac{-2\langle\log h\rangle-\log\left(\langle h^{2}\rangle_{Q}\cdot K(N)\right)+\log(N\beta\langle h^{2}\rangle)+\log\log\left(\frac{N\langle h^{2}\rangle}{\beta}\right)-\gamma}{\log N+\log\langle h^{2}\rangle-\log\beta}+\frac{1}{2}$
2.3. Path lengths via Landmarks in random networks
@@
The proof of Theorem 1 is in Appendix A.1. In (1), the design parameters are $\langle h^{2}\rangle_{Q}$ and $K(N)$ which are related to what kind of landmarks are chosen, and how many of them, respectively. Next, we bound the average of the minimum path length among the detours via landmarks. In what follows, we will assume $K(N)=o(N)$ and $K(N)\to\infty$ as $N\to\infty$, implying that the number of landmarks is chosen to be not too large compared to $N$.
@@
of the shortest detour via landmarks, denoted by $\bar{l}$, is bounded above as (2) where $\gamma\approx 0.5772$ is the Euler’s constant.
The proof of Theorem 2 is in Appendix A.2. We apply the results to some well-known models for random graphs.
@@ 2.4. Erdős-Rényi Model
The Erdős-Rényi (ER) model (Erdős et al., 1960) is a classical random graph in which every node pair is connected with a common probability. For ER graphs, model parameters $\beta$ and $h$ can be set as $\beta=\langle k\rangle N$ and $h\equiv\langle k\rangle$ respectively, where $\langle k\rangle$ denotes the mean degree of nodes (Boguná and Pastor-Satorras, 2003). From (3), we have $q_{ij}=\langle k\rangle/N$, i.e., the probability of edge formation between any node pair is constant. The node degree follows the Poisson distribution with mean $\langle k\rangle$ for large $N$. Assume $\langle k\rangle$ is a finite constant. By applying the model parameters to (2), the average distance of the shortest detour via landmarks in ER network, denoted by $\bar{l}_{\textrm{ER}}$, is bounded above as
@@
asymptotically in $N$. The average length of the shortest paths in ER graphs without landmarks, denoted by $\bar{l}^{*}_{\textrm{ER}}$, is given by (Boguná and Pastor-Satorras, 2003; Fronczak et al., 2004)
@@
By comparing (5) and (6), we observe that the detour via landmarks incurs the overhead of at most factor 2. This is because nodes in ER graphs appear homogeneous, and thus the path length to and from landmarks are on average similar to the inter-node distance. Thus the average distance of a detour will be twice the direct distance. However, (5) implies that the minimum detour distance can be reduced by using multiple $(K(N)>1)$ landmarks. The reduction can be substantial, e.g., if $K(N)=N^{1-\varepsilon}$ for some $\varepsilon\in(0,1)$, we have, from (5),
@@ 2.5. Barabási-Albert Model
The Barabási-Albert (BA) model (Barabási and Albert, 1999) generates random graphs with preferential attachment. BA networks are characterized by continuous growth over time: initially there are $m$ nodes, and new nodes arrive to the network over time. Due to preferential attachment, the probability of the connection of the existing node to the newly arriving node is proportional to its degree. The probability of an edge in BA networks is shown to be (Boguná and Pastor-Satorras, 2003)
@@
with $h_{i}=1/\sqrt{t_{i}}$ and $\beta=\frac{m}{2}$, where $t_{i}$ is the time of arrival of node $i$. Since the probability of a newly arriving node connecting to node $i$ is proportional to $h_{i}$, the degree of nodes with large $h_{i}$ is likely to be high. For large $N$, the distribution of $h$ is derived as (Fronczak et al., 2004)
@@
By applying $\rho(\cdot)$ to (2), we bound the average path length with landmarks denoted by $\bar{l}_{\textrm{BA}}$ as (4). Unlike ER graphs, there exists a landmark selection strategy which achieves the asymptotically optimal distance, despite using a small number of landmarks relative to the network size.
@@
[…]
asymptotically in $N$.
The proof of Theorem 3 is in Appendix A.3. We claim the optimality of the strategy in Theorem 3 in the following sense. Since $K(N)$ is slowly increasing in $N$, e.g., $K(N)=O(\log N)$, a feasible strategy for Theorem 3 is to choose $O(\log N)$ landmarks from top–$O(\log^{2}N)$ degree nodes. For such $K(N)$, the numerator of the RHS of (8) is $\approx\log N$ for large $N$. Meanwhile, the average length of shortest paths in BA networks is given by (Cohen and Havlin, 2003)
$\bar{l}^{*}_{\textrm{BA}}=\dfrac{\log N}{\log\log N}$
Thus, we conclude that $\bar{l}_{\textrm{BA}}$ is asymptotically equal to $\bar{l}^{*}_{\textrm{BA}}$. This implies that the shortest detour via landmarks under our strategy has the same length as the shortest path on average in the asymptotic sense.
Verifying Analysis. Numerical verification of the bounds derived in this section is provided in Appendix B.
Figure 1. Overview of HPLC. 1⃝ Partition the graph into $K$ clusters using FluidC, select landmarks based on degrees, and compute distance vectors of nodes. 2⃝ Construct a landmark graph to compute membership vectors based on eigenvectors of graph Laplacian. 3⃝ Compute positional embeddings by combining membership and distance vectors and passing them through an encoder. 4⃝ Concatenate positional embeddings and node features, and project them onto cluster-group embedding spaces. 5⃝ Neighborhood aggregation using GNNs. $\oplus$ denotes concatenation.

### v2

@@ Abstract.
Learning positional information of nodes in a graph is important for link prediction tasks. We propose a representation of positional information using representative nodes called landmarks. A small number of nodes with high degree centrality are selected as landmarks which serve as reference points for the nodes’ positions. We justify this selection strategy for well-known random graph models and derive closed-form bounds on the average path lengths involving landmarks. In a model for power-law graphs, we prove that landmarks provide asymptotically exact information on inter-node distances. We apply theoretical insights to real-world graphs and propose Hierarchical Position embedding with Landmarks and Clustering (HPLC). HPLC combines landmark selection and graph clustering, i.e., the graph is partitioned into densely connected clusters in which nodes with the highest degree are selected as landmarks. HPLC leverages the positional information of nodes based on landmarks at various levels of hierarchy such as nodes’ distances to landmarks, inter-landmark distances and hierarchical grouping of clusters. Experiments show that HPLC achieves state-of-the-art performances of link prediction on various datasets in terms of HIT@K, MRR, and AUC. The code is available at https://github.com/kmswin1/HPLC.
@@ 1. Introduction
Message passing GNNs (MPGNNs) (Gilmer et al., 2017; Kipf and Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018) have been successful in learning structural node representations through neighborhood aggregation. However, standard MPGNNs achieved subpar performances on link prediction tasks due to their inability to distinguish isomorphic nodes. In order to tackle isomorphism, subgraph-based methods (Zhang and Chen, 2018; Pan et al., 2022) were proposed for link prediction based on dynamic node embeddings from enclosing subgraphs. Labeling methods (Li et al., 2020; Zhang et al., 2021) assign node labels to identify each node via hashing or aggregation functions. An important line of research is the position-based methods  (You et al., 2019; Dwivedi et al., 2020; Kreuzer et al., 2021; Wang et al., 2022) which propose to learn and utilize the positional information of nodes.
@@
The positional information of nodes is useful for link prediction because, e.g., isomorphic nodes can be distinguished by their positions, and the link formation between a node pair may be related to their relative distance (Srinivasan and Ribeiro, 2020). The position of a node can be defined as its distances relative to other nodes. A relevant result by Bourgain (Bourgain, 1985) states that inter-node distances can be embedded into Euclidean space with small errors. Linial (Linial et al., 1995) and P-GNN (You et al., 2019) focus on constructing Bourgain’s embeddings based on the nodes’ distances to random node sets. Another line of positional encoding used the eigenvectors of graph Laplacian (Dwivedi and Bresson, 2021; Kreuzer et al., 2021) based on spectral graph theory. However, the aforementioned methods have convergence issues and may not scale well for large or dense graphs (Wang et al., 2022). For example, Laplacian methods have stability issues (Wang et al., 2022) and may not outperform methods based on structural features (Zhang and Chen, 2018). Thus, there is significant room for improving both performance and scalability of positional node embedding for link prediction.
@@
In this paper, we propose an effective and efficient representation of the positional information of nodes. We select a small number of representative nodes called landmarks and impose hierarchy on the graph by associating each node with a landmark. Each node computes distances to the landmarks, and each landmark computes distances to other landmarks. Such distance information is combined so as to represent nodes’ positional information. Importantly, we select landmarks and organize the graph in a principled way, unlike previous methods using random selection, e.g., (Bourgain, 1985; Linial et al., 1995; You et al., 2019).
@@
The key question is how to select landmarks. We advocate the selection based on degree centrality which is motivated from the theory of network science. In network models with preferential attachment (PA), nodes with high degrees called hubs are likely to emerge (Barabási and Albert, 1999). PA is a process such that, if a new node joins the graph, it is more likely to connect to nodes with higher degrees. PA induces the power-law distribution of node degrees, and the network exhibits scale-free property (Barabási and Albert, 1999). Hubs are abundant in social/citation networks and World Wide Web, and play a central role in characterizing the network.
@@
Motivated by the theory, we propose Hierarchical Position embedding with Landmarks and Clustering (HPLC). HPLC partitions a graph into $O(\log N)$ clusters which are locally dense, and appoints the node with the highest degree in the cluster as the landmark. Our intention is to bridge gap between theory and practice: hubs may not be present in real-world graphs. Thus, it is important to distribute the landmarks evenly over the graph through clustering, so that nodes can access nearby/local landmarks. Next, we form a graph of higher hierarchy, i.e., the graph of landmarks, and compute its Laplacian encoding based on the inter-landmark distances. The encoding is assigned as a membership to the nodes belonging to the cluster so as to learn positional features at the cluster level. We further optimize our model using the encoding based on hierarchical grouping of clusters. The computation of HPLC can be mainly done during preprocessing, incurring low computational costs. Our experiments on 7 datasets with 16 baseline methods show that HPLC achieves state-of-the-art performances, demonstrating its effectiveness especially over prior position-/distance-based methods.
Our contributions are summarized as follows: 1) we propose HPLC, a highly efficient algorithm for link prediction using hierarchical position embedding based on landmarks combined with graph clustering; 2) building upon network science, we derive closed-form bounds on average path lengths via detours for well-known random graphs which, to our belief, are important theoretical findings; 3) we conduct extensive experiments to show that HPLC achieves state-of-the-art performances in most cases.
@@ 2.1. Notation
We consider undirected graph $G=(V,E)$ where $V$ and $E\subseteq V\times V$ denote the set of vertices and edges, respectively. Let $N$ denote the number of nodes, or $N=|V|$. $\mathbf{A}\in\mathbb{R}^{N\times N}$ denotes the adjacency matrix of $G$. $d(v,u)$ denotes the geodesic (shortest-path) distance between $v$ and $u$. Node attributes are defined as $X=\{x_{1},...,x_{N}\}$ where $x_{i}\in\mathbb{R}^{n}$ denotes the feature vector of node $i$. We consider methods of embedding $X$ into the latent space as vectors $Z=\{z_{1},...,z_{N}\},z_{i}\in\mathbb{R}^{m}$. We study the node-pair-level task of estimating the link probability between $u$ and $v$ from $z_{u}$ and $z_{v}$.
@@ 2.2. Representation of Positions using Landmarks
$D(v)$ can be used to represent the (approximate) position of node $v$. If the positions of $u$ and $v$ are to be used for link prediction, one can combine $D(u)$ and $D(v)$ (or the embeddings thereof) to estimate the probability of link formation.
@@
We examine the achievable accuracy of distance estimates based on the detours via landmarks. Previous works on distance estimates using auxiliary nodes relied on heuristics (Potamias et al., 2009) or complex algorithms for arbitrary graphs (Linial et al., 1995; You et al., 2019). However, real-world graphs have structure, e.g., preferential attachment or certain distribution of node degrees. Presumably, one can exploit such structure to devise efficient algorithms. The key design questions are: how to select good landmarks, and how many of them? Drawing upon the theory of network science, we analyze a well-known class of random graphs, derive the detour distances via landmarks, and glean design insights from the analysis.
(1) $P(L_{ij}>s)=\exp\bigg{[}-\frac{h_{i}h_{j}}{\beta N\langle h^{2}\rangle}\cdot\langle h^{2}\rangle_{Q}\cdot K(N)\cdot(s-1)\bigg{(}\frac{\langle h^{2}\rangle N}{\beta}\bigg{)}^{s-1}\bigg{]},\quad s=1,2,\cdots$
(2) $\bar{l}\leq\dfrac{-2\langle\log h\rangle-\log\left(\langle h^{2}\rangle_{Q}\cdot K(N)\right)+\log(N\beta\langle h^{2}\rangle)+\log\log\left(\frac{N\langle h^{2}\rangle}{\beta}\right)-\gamma}{\log N+\log\langle h^{2}\rangle-\log\beta}+\frac{1}{2}$
2.3. Path lengths via Landmarks in random graphs
@@
The proof of Theorem 1 is in Appendix A.1. In (1), the design parameters are $\langle h^{2}\rangle_{Q}$ and $K(N)$ which are related to what kind of landmarks are chosen, and how many of them, respectively. Next, we bound the average distance of the shortest detour via landmarks. In what follows, we will assume $K(N)=o(N)$ and $K(N)\to\infty$ as $N\to\infty$, implying that the number of landmarks is chosen to be not too large compared to $N$.
@@
The average distance of the shortest detour via landmarks, denoted by $\bar{l}$, is bounded above as (2) where $\gamma\approx 0.5772$ is the Euler’s constant.
The proof of Theorem 2 is in Appendix A.2. Next, we apply the results to some well-known models for random graphs.
@@ 2.4. Erdős-Rényi Model
The Erdős-Rényi (ER) model (Erdős et al., 1960) is a classical random graph in which every node pair is connected with a common probability. For ER graphs, model parameters $\beta$ and $h$ can be set as $\beta=\langle k\rangle N$ and $h\equiv\langle k\rangle$ respectively, where $\langle k\rangle$ denotes the mean degree of nodes (Boguná and Pastor-Satorras, 2003). From (3), we have $q_{ij}=\langle k\rangle/N$, i.e., the probability of edge formation between any node pair is constant. The node degree follows the Poisson distribution with mean $\langle k\rangle$ for large $N$.
We apply our analysis to ER graphs as follows. By plugging in the model parameters to (2), the average distance of the shortest detour via landmarks in ER graphs, denoted by $\bar{l}_{\textrm{ER}}$, is bounded above as
@@
asymptotically in $N$. Meanwhile, the average inter-node distance in ER graphs denoted by $\bar{l}^{*}_{\textrm{ER}}$, is given by (Boguná and Pastor-Satorras, 2003; Fronczak et al., 2004)
@@
By comparing (5) and (6), we observe that the detour via landmarks incurs the overhead of at most factor 2. This is because, the nodes in ER graphs appear homogeneous, and thus the path length to and from landmarks are on average similar to the inter-node distance. Thus, the average distance of a detour will be twice the direct distance. However, (5) implies that the minimum detour distance can be reduced by using multiple $(K(N)>1)$ landmarks. The reduction can be substantial, e.g., if $K(N)=N^{1-\varepsilon}$ for some $\varepsilon\in(0,1)$, we have, from (5),
@@ 2.5. Barabási-Albert Model
The Barabási-Albert (BA) model (Barabási and Albert, 1999) generates random graphs with preferential attachment. BA graphs are characterized by continuous growth over time: initially there are $m$ nodes, and new nodes arrive to the network over time. Due to preferential attachment, the probability of a newly arriving node connecting to the existing node is proportional to its degree. The probability of an edge in BA graphs is shown to be (Boguná and Pastor-Satorras, 2003)
@@
with $h_{i}=1/\sqrt{t_{i}}$ and $\beta=\frac{2}{m}$, where $t_{i}$ is the time of arrival of node $i$. Since the probability of a newly arriving node connecting to node $i$ is proportional to $h_{i}$, the degree of nodes with large $h_{i}$ is likely to be high. For large $N$, the distribution of $h$ is derived as (Fronczak et al., 2004)
@@
By applying $\rho(\cdot)$ to (2), we bound the average path length with landmarks denoted by $\bar{l}_{\textrm{BA}}$ as (4). Importantly, unlike ER graphs, there exists a landmark selection strategy which achieves the asymptotically optimal distance, despite using a small number of landmarks relative to the network size.
@@
Suppose $K(N)$ landmarks are randomly selected from the nodes with top-$(\log N)\cdot K(N)$ largest values of $h$. Assume $m=O(1)$. The average distance of the shortest detour via landmarks, denoted by $\bar{l}_{\textrm{BA}}$, is bounded as
(8) $\bar{l}_{\textrm{BA}}\leq\dfrac{\log N-\log\log K(N)+2\log\log N}{\log\log N}$
asymptotically in $N$.
The proof of Theorem 3 is in Appendix A.3. We claim the optimality of the strategy in Theorem 3 in the following sense. Since $K(N)$ is slowly increasing in $N$, e.g., $K(N)=O(\log N)$, a feasible strategy for Theorem 3 is to choose $O(\log N)$ landmarks from top–$O(\log^{2}N)$ degree nodes. For such $K(N)$, the numerator of the RHS of (8) is $\approx\log N$ for large $N$. Meanwhile, the average length of shortest paths in BA graphs is given by (Cohen and Havlin, 2003)
(9) $\bar{l}^{*}_{\textrm{BA}}=\dfrac{\log N}{\log\log N}$
By comparing (8) and (9), we conclude that $\bar{l}_{\textrm{BA}}$ is asymptotically equal to $\bar{l}^{*}_{\textrm{BA}}$. This implies that the shortest detour via landmarks under our strategy has on average the same length as the shortest path in the asymptotic sense.
Summary of Analysis. ER and BA models represent two contrasting cases of degree distributions. The degree distribution of ER graph is highly concentrated, i.e., the variation of node degrees is small, or hubs are absent. By contrast, the degree of BA graphs follows power-law distribution, i.e., the variation of node degree is large, and hubs are present. Our analysis shows that $(1+\epsilon)$-factor approximation (ER) and asymptotic optimality (BA) are achievable by detour via landmarks. The derived bounds are numerically verified via simulation in Appendix B.